X1 X2 Y X1 1.000 0.454 0.585 X2 0.454 1.000 0.342 Y 0.585 0.342 1.000 n=29...

X1 X2 Y
X1 1.000 0.454 0.585
X2 0.454 1.000 0.342
Y 0.585 0.342 1.000

n=29

Calculated the adjusted R2
Answer #1
Correlation Matrix
X1 X2 Y
X1 1.00 0.45 0.59
X2 0.45 1.00 0.34
Y 0.59 0.34 1.00

rx1 = correlation between Y and X1 = 0.59

rx2 = correlation between Y and X2 = 0.34

r12 = correlation between X2 and X1 = 0.45

R2 is given by, R2 = ( rx12 + rx22 - 2rx1rx2r12 ) / ( 1 - r122 )

= (0.592 + 0.342 - 2 *0.59*0.34*0.45) / ( 1 - 0.452 )

= 0.2775/ 0.7938 = 0.35

n = no of observations = 29

k = no of independents used = 2

Adjusted R2 = 1 - (1 - R2 )*[(n-1)/(n-(k+1))]

= 1 - (1 - 0.35)*[(29-1)/(29-(2+1))]

= 1 - (1 - 0.35)*[28/26]

= 1 - 0.7 = 0.3